OpenStudy (anonymous):

If the length of a rectangle is decreased by 4cm and the width is increased by 5 cm, the result will be a square, the area of which will be 40 cm^2 greater than the area of the rectangle. Find the area of the rectangle.

OpenStudy (anonymous):

360

OpenStudy (anonymous):

 (x + 4)(x - 5) = x^2 - 40 use algebra to find x = 20.. so, area of square is 400 and area of rectangle is 360!